Session: Feb 2011 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Pablo de Vicente at Fri Mar 04 10:13:32 2011) Minor failures: WIND
The observation completed but we had snow and part of the lower panels were covered with snow giving high Tsys during some intervals of the observation.

Medicina (roma at Fri Mar 04 10:15:17 2011) Success: RFI
All scan recorded From 08:03 UT some problems for RFI cloudy sky

Torun (K. Borkowski at Fri Mar 04 10:58:03 2011) Success:
No known problems.

Westerbork (Geert Kuper at Fri Mar 04 12:52:27 2011) Minor failures:
We had an error condition in the input section of TADUMax until 07:53 UT , but the data could still be ok. We tried to reset the Tadu 3 times. The last one cleared the error message. 03:11-03:13 UT no data (restart observation) 03:20-03:24 UT no data (restart observation) 07:53-08:12 UT no data (backend reboot)

Onsala (Michael Lindqvist at Fri Mar 04 20:32:46 2011) Success:
No known problems.

Shanghai (BO XIA at Mon Mar 07 10:17:56 2011) Success:

Hartebeesthoek (Jonathan Quick at Tue Mar 08 13:06:25 2011) Success:
IFD comms errors, no other known problems

Effelsberg (Uwe Bach at Wed Mar 09 14:35:56 2011) Success:

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:31:13 2011) Success:
Receiver was replaced prior to this experiment. No known problems.